# KVCrush: Key Value Cache size-reduction using similarity in head-behaviour

### v1

@@ IV-A3 Models
KVCrush* | 25.62 | 29.48 | 39.56 | 44.05 | 36.20 | 21.62 | 26.24 | 23.28 | 26.52 | 74.00 | 90.56 | 42.24 | 5.83 | 69.50 | 54.49 | 40.61 |
IV-B Tuning KVCrush for accuracy and efficiency (RQ1)
IV-B1 Selecting Anchor Points in KVCrush
KVCrush employs an anchor point within the same binary space (a binary vector with a length equal to the number of heads) to form clusters of tokens based on their hamming distance to this anchor. We present our experiments with anchor points in Table II. KVCrush outperforms the baseline H2O even using generic anchor points like random, mean and alternate 0s and 1s.
IV-B2 Cache budget partitioning
As discussed in Section III, the cache budget allocated for the KV Cache is divided into two distinct segments: , which is utilized by the baseline method to identify tokens with high attention weights, and , which is employed by KVCrush to select representative tokens.
Figure 5 illustrates the impact of varying the percentage of the budget allocated to KVCrush on overall accuracy. For certain workloads, such as narrativeqa, the optimal cache budget can reach up to of the total budget. In contrast, for other workloads, the optimal budget typically ranges between and of the available budget. In Section IV-F, we compared our method with baseline approaches using a static cache budget partitioning scheme, allocating a fixed of the total budget to KVCrush. This approach could potentially be enhanced by dynamically partitioning the budget based on attention weights distribution, which we plan to explore in future work.
IV-C KMeans vs KVCrush (RQ2)
Traditional clustering algorithms exhibit significant inefficiencies in KV Cache compression due to two primary factors. Firstly, the large size of input vectors poses a challenge. For instance, in a typical Llama3-8B model, each input key (or value) vector comprises 128 FP16 values. By employing kvcrush.representation, a binary vector of length 32 (corresponding to the number of heads) is generated to represent each token (key-value pair). This approach not only reduces the data size by a factor of 64 but also enables faster distance computations using Hamming distances. Secondly, the time complexity associated with clustering algorithms for data grouping is substantial. Utilizing kvcrush.grouping, the selection of values to be retained in the cache is achieved in time, in contrast to the time complexity of k-means clustering.
In Figure 6, we compare KVCrush and KMeans in terms of accuracy and inference latency using the GSM8K dataset. While KMeans achieves slightly higher accuracy than KVCrush (Figure 6(a)), it incurs approximately additional overhead in the inference latency pipeline. In contrast, KVCrush provides a reasonable accuracy improvement over the pure H2O baseline, with an insignificant overhead of less than .
IV-D Accuracy implications of integrating KVCrush with other methods(RQ3)
We assess the impact on accuracy when integrating KVCrush with other methods, utilizing cumulative attention-weights as the importance metric. As illustrated in Figure 2, KVCrush can be seamlessly integrated with these methods, which prioritize selecting important tokens, while KVCrush emphasizes selecting representative tokens based on the head behavior of these tokens (or pages). Figure 7 demonstrates the accuracy improvement provided by KVCrush to the baseline methods for the 2wikimqa dataset. Figure 7(b) demonstrates that KVCrush can enhance the accuracy of baseline methods even when it evicts chunks instead of individual tokens.
IV-E Evalution of KVCrush in Paged KV settings (RQ4)
To evaluate KVCrush in paged KV settings, we utilized an H2O baseline that aggregates the row-wise sum of attention weights at the page level to evict low-importance pages. For KVCrush, the binary vector of a page is formed by concatenating the binary head vectors of all tokens within that page. Figure 8 shows that Paged-KVCrush outperforms paged-H2O on both GSM8K and XSUM datasets.
IV-F Accuracy vs Baselines (RQ5)
The accuracy of KVCrush, in comparison with baseline methods, is detailed in Table III. For this evaluation, a cache budget of 2048 was used. KVCrush employs a static cache budget partitioning scheme, wherein of the total budget (i.e., 512) is allocated to KVCrush for selecting representative tokens. The remaining is distributed to the accompanying method (H2O, SnapKV, or PyramidKV) for selecting high-attention tokens. The key insights are as follows:
@@ IV-F Accuracy vs Baselines (RQ5)
V Conclusion and Future Work

### v2

@@ 4.2.1 Accuracy
| PyramidKV+KVCrush | 25.57 | 29.48 | 38.97 | 44.03 | 35.75 | 21.62 | 26.21 | 23.28 | 26.52 | 74.00 | 90.56 | 42.24 | 5.17 | 69.50 | 55.49 | 40.56 |
| KVCrush* | 25.62 | 29.48 | 39.56 | 44.05 | 36.20 | 21.62 | 26.24 | 23.28 | 26.52 | 74.00 | 90.56 | 42.24 | 5.83 | 69.50 | 55.49 | 40.68 |
The accuracy of KVCrush, in comparison with baseline methods, is detailed in Table 3. For this evaluation, a cache budget of 2048 was used. KVCrush employs a static cache budget partitioning scheme, wherein of the total budget (i.e., 512) is allocated to KVCrush for selecting representative tokens. The remaining is distributed to the accompanying method (H2O, SnapKV, or PyramidKV) for selecting high-attention tokens. The key insights are as follows:
@@
4.2.2 Inference Latency Overhead
fig:latency_breakdown
We evaluate the runtime latency of different KV cache strategies using a LongBench microbenchmark. The experiments were conducted on a system equipped with an Intel® Xeon® Platinum 8470 processor, simulating realistic CPU-bound server inference settings. The cache budget and associated compression parameters are detailed in Table 1.
Figure LABEL:fig:latency_breakdown shows the normalized latency breakdown across three components: memory access, compute, and pruning overhead. We observe that while the full cache baseline incurs the highest memory access latency, both H2O and H2O+KVCrush significantly reduce this cost by reducing the KV cache size by . Importantly, the additional pruning overhead introduced by KVCrush remains minimal—only higher than H2O—while achieving higher accuracy. This demonstrates that KVCrush can be deployed with negligible runtime cost while improving performance.
4.3 Ablation Study
4.3.1 Selecting Anchor Points
KVCrush employs an anchor point within the same binary space (a binary vector with a length equal to the number of heads) to form clusters of tokens based on their hamming distance to this anchor. As shown in Table 2, KVCrush outperforms the H2O baseline across all anchor-point strategies, including simple choices such as random, mean, and alternating 0–1 vectors. The 512-token budget shown is illustrative; similar trends are observed for other cache budgets.
4.3.2 Cache budget partitioning
As discussed in Section 3, the cache budget allocated for the KV Cache is divided into two distinct segments: , utilized by the baseline method to identify tokens with high attention weights, and , used by KVCrush to select representative tokens.
Figure 5 illustrates the impact of varying the percentage of the budget allocated to KVCrush on overall accuracy. For certain workloads, such as narrativeqa, the optimal cache budget can reach up to of the total budget. In contrast, for other workloads, the optimal budget typically ranges between and of the available budget. In Section 4.2.1, we compared our method with baseline approaches using a static cache budget partitioning scheme, allocating a fixed of the total budget to KVCrush. This approach could potentially be enhanced by dynamically partitioning the budget based on attention weights distribution, which we plan to explore in future work.
fig:kmeans
fig:kvcrush_int
4.3.3 KMeans vs KVCrush
Traditional clustering algorithms exhibit inefficiencies in KV Cache compression due to two primary factors. Firstly, the large size of input vectors poses a challenge. For instance, in Llama3-8B model, each input key (or value) vector comprises 128 FP16 values. By employing kvcrush.representation, a binary vector of length 32 (corresponding to the number of heads) is generated to represent each token (key-value pair). This approach not only reduces the data size by a factor of 64 but also enables faster distance computations using Hamming distances. Secondly, the time complexity for clustering algorithms is substantial. Utilizing kvcrush.grouping, the selection of values to be retained in the cache is achieved in time, in contrast to the time complexity of k-means clustering.
In Figure LABEL:fig:kmeans, we compare KVCrush and KMeans in terms of accuracy and inference latency using the GSM8K dataset. While KMeans achieves slightly higher accuracy than KVCrush (Figure LABEL:fig:kmeans_acc), it incurs approximately additional overhead in the inference latency pipeline. In contrast, KVCrush provides a reasonable accuracy improvement over the pure H2O baseline, with an insignificant overhead of less than .
4.3.4 Accuracy implications of integrating KVCrush with other methods
We assess the impact on accuracy when integrating KVCrush with other methods, utilizing cumulative attention-weights as the importance metric. Figure LABEL:fig:kvcrush_int demonstrates the accuracy improvement provided by KVCrush to the baseline methods for the 2wikimqa dataset. These baseline methods will prioritize selecting important tokens. As illustrated in Figure 1, KVCrush not only seamlessly integrates with these, but also improves upon their accuracies by selecting representative tokens (or pages) based on the head behavior. As shown in figure, KVCrush also enhances the accuracy of the baseline methods even when they operate at chunks instead of individual tokens.
fig:kvcrush_paged
4.3.5 Evaluation of KVCrush in Paged KV settings
To evaluate KVCrush in paged KV settings, we utilized an H2O baseline that aggregates the row-wise sum of attention weights at the page level to evict low-importance pages. For KVCrush, the binary vector of a page is formed by concatenating the binary head vectors of all tokens within that page. Figure LABEL:fig:kvcrush_paged shows that Paged-KVCrush outperforms paged-H2O on both GSM8K and XSUM datasets.
5 Conclusion and Future Work